### v1

# 3HDM at the ILC

Abstract
The Three-Higgs-Doublet Model (3HDM) extends the Standard Model by introducing two additional scalar doublets, leading to a rich spectrum of new particles: three neutral CP-even Higgs bosons ($h_{1}$, $H_{2}$, $H_{3}$), two neutral CP-odd Higgs bosons ($A_{2}$, $A_{3}$), and two charged Higgs bosons ($H_{2}^{+}$, $H_{3}^{+}$). In this work, we present a phenomenological study of the 3HDM at the future International Linear Collider (ILC) with a center-of-mass energy of $\sqrt{s}=1000\,\text{GeV}$. Applying a comprehensive set of theoretical and experimental constraints, we identify promising new physics signals with sufficiently large production cross-sections. Our analysis shows that $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $A_{2}A_{2}Z$, $H_{2}H_{2}^{\pm}W^{\mp}$, $A_{2}H_{2}^{\pm}W^{\mp}$ and $H_{1}H_{2}A_{2}$ are among the most sensitive channels to probe this extended Higgs sector. We demonstrate that a future ILC would offer a powerful platform to test these interactions and discover these heavier Higgs bosons thus providing evidence of physics beyond the Standard Model.
I Introduction
The Standard Model (SM) of particle physics has been the most successful theory so far, verified and established by past and current experimental setups. The final validation of the SM came with the discovery of the Higgs boson at the Large Hadron Collider (LHC) in 2012 [1, 2]. Although the SM has been the cornerstone of particle physics, it has long been known that it needs to be supplanted with new dynamics at the TeV scale owing to many theoretical and experimental reasons - the evidence of dark matter and neutrino mass are two of the most pressing theoretical issues that do not find a resolution within the SM. One of the many ways the theoretical framework of the SM can be extended is by adding additional scalar sectors, such as singlet [3, 4] and doublet extensions [5], and their generalization to multi-Higgs doublet models [6, 7]. Of late, along these lines, the 3HDM has emerged as a potentially interesting and promising source of investigation [8, 9]. These models have been key in addressing unresolved questions, such as providing new sources of CP violation which are crucial for explaining baryogenesis [10], enabling strong first-order electroweak phase transitions, a necessary condition for successful electroweak baryogenesis [11]. In certain realizations, the extended scalar sector can also offer viable dark matter candidates through neutral scalars [12]. Beyond cosmological implications, such models predict a richer Higgs spectrum with distinct collider signatures, offering promising discovery opportunities at current and future colliders like the LHC and ILC. While models with two Higgs doublets have been extensively studied [5], the 3HDM offers an even richer framework for addressing fundamental questions beyond the Standard Model. Furthermore, the additional scalar degrees of freedom enable more flexible Yukawa structures, providing potential explanations for the observed patterns of fermion masses and mixing [13]. From a phenomenological perspective, the extended scalar spectrum enables a variety of novel collider signatures including multiple charged and neutral Higgs bosons that can be probed at future experiments such as the International Linear Collider (ILC) or the Future Circular Collider (FCC-ee).
In this paper, we explore the phenomenological aspects of the 3HDM from the standpoint of the future ILC experiment. As the ILC global project proposal [14, 15] continues to evolve and gain traction within the community, it is timely and important to investigate the avenues through which various Beyond Standard Model (BSM) scenarios could manifest in its setup. With this motivation, we investigate the phenomenology of the heavy Higgs bosons in the 3HDM at the ILC operating at a center-of-mass energy of $\sqrt{s}=$ 1000 GeV [16]. We focus on channels that are expected to have reasonable cross sections while operating with benchmark points allowed by all existing theoretical and experimental constraints. The paper is organized as follows: in Sec. II, we provide a brief overview of the 3HDM concentraing on the scalar structure, and list the theoretical and experimental constraints on the model. The phenomenological analysis of the various relevant channels at the ILC is presented in Sec. III, and we present our conclusions in Sec. IV.
II Model and Constraints
In this section, we present a concise overview of the model, focusing on the scalar and Yukawa sectors. We outline the structure of the scalar potential, detail the transformations between gauge and mass eigenstates, and comment on the Yukawa sector. We then discuss the alignment limit conditions, followed by a summary of the theoretical and experimental constraints imposed to identify viable benchmark points for the subsequent phenomenological analysis.
II.1 Scalar Sector
The Three Higgs Doublet model extends SM scalar sector by introducing two additional $SU(2)_{L}$ doublets represented as
$\Phi_{k}=\begin{pmatrix}\phi_{k}^{+}\\ \frac{v_{k}+p_{k}+in_{k}}{\sqrt{2}}\end{pmatrix},$ (1)
where $k=1,2,3$. We consider a $Z_{3}$ symmetric potential under which the Higgs fields transform as
$\Phi_{1}\rightarrow\omega\Phi_{1},\;\;\Phi_{2}\rightarrow\omega^{2}\Phi_{2},\;\textrm{and}\;\Phi_{3}\rightarrow\Phi_{3},$ (2)
where $\omega=e^{2\pi i/3}$ are the cube roots of unity. The most general $SU(2)_{L}\times U(1)_{Y}$ invariant potential preserving the $Z_{3}$ symmetry is given by
$\begin{split}V&=m_{11}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+m_{22}^{2}(\Phi_{2}^{\dagger}\Phi_{2})+m_{33}^{2}(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{1}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\lambda_{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}+\lambda_{3}(\Phi_{3}^{\dagger}\Phi_{3})^{2}\\ &+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\lambda_{5}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{3})+\lambda_{6}(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{7}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})+\lambda_{8}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{1})+\lambda_{9}(\Phi_{2}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{2})\\ &+[\lambda_{10}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{3})+\lambda_{11}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{2})+\lambda_{12}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{2}^{\dagger}\Phi_{3})+h.c.].\\ &\end{split}$ (3)
Here, in general $\lambda_{1,...,9}$ are real parameters, while $\lambda_{10},\lambda_{11},\lambda_{12}$ can be complex. However, here we disregard the complex phases of the potential to prevent mixing between the CP-odd and CP-even states [8]. The relevant mass terms for the CP-even Higgs are deduced from the scalar potential in Eqn. 3, and can be written as
$V_{p}^{mass}\supset\begin{pmatrix}p_{1}&p_{2}&p_{3}\end{pmatrix}\frac{\mathcal{M}^{2}_{S}}{2}\begin{pmatrix}p_{1}\\ p_{2}\\ p_{3}\end{pmatrix},$
As a real symmetric matrix, it can be diagonalized through an orthogonal transformation characterized by the matrix $O_{\alpha}$. The diagonalization condition is given by
$O_{\alpha}.\mathcal{M}^{2}_{S}.O_{\alpha}^{T}=\begin{pmatrix}m_{H1}^{2}&0&0\\ 0&m_{H2}^{2}&0\\ 0&0&m_{H3}^{2}\end{pmatrix}$ (4)
The explicit relations of $\lambda_{1,2,..,6}$ can be expressed in terms of the masses of the three CP-even Higgs bosons and three mixing angles. The gauge eigenstates can be related to the mass eigenstates thus:
$\begin{split}H_{1}&=c_{\alpha_{2}}c_{\alpha_{1}}p_{1}+c_{\alpha_{2}}s_{\alpha_{1}}p_{2}+s_{\alpha_{2}}p_{3},\\ H_{2}&=-(c_{\alpha_{3}}s_{\alpha_{1}}+s_{\alpha_{3}}s_{\alpha_{2}}c_{\alpha_{1}})p_{1}+(c_{\alpha_{3}}c_{\alpha_{1}}-s_{\alpha_{3}}s_{\alpha_{2}}s_{\alpha_{1}})p_{2}+(s_{\alpha_{3}}c_{\alpha_{2}})p_{3},\,\textrm{and}\\ H_{3}&=(s_{\alpha_{3}}s_{\alpha_{1}}-c_{\alpha_{3}}s_{\alpha_{2}}c_{\alpha_{1}})p_{1}-(s_{\alpha_{3}}c_{\alpha_{1}}+c_{\alpha_{3}}s_{\alpha_{2}}s_{\alpha_{1}})p_{2}+(c_{\alpha_{3}}c_{\alpha_{2}})p_{3}.\end{split}$ (5)
The mass terms for the charged Higgses can similarly be extracted from the scalar potential given in Eqn. 3:
$V_{C}^{mass}\supset\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\mathcal{M}^{2}_{\phi^{\pm}}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix},$
The diagonalization of this matrix is performed through two successive transformations, and the relation between the gauge and mass eigenstates is given by
$\begin{pmatrix}\phi_{1}^{\pm}\\ \phi_{2}^{\pm}\\ \phi_{3}^{\pm}\end{pmatrix}=(O_{\gamma 2}.O_{\beta})^{\dagger}\begin{pmatrix}G^{\pm}\\ H_{2}^{\pm}\\ H_{3}^{\pm}\end{pmatrix},$
where
$(O_{\gamma 2}.O_{\beta})^{\dagger}=\begin{pmatrix}c_{\beta 1}c_{\beta 2}&-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2}\\ c_{\beta 2}s_{\beta 1}&c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2}\\ s_{\beta 2}&-c_{\beta 2}s_{\gamma 2}&c_{\beta 2}c_{\gamma 2}\end{pmatrix}.$ (6)
The gauge eigenstates can be represented in terms of mass eigenstates as
$\begin{split}G^{\pm}&=c_{\beta 1}c_{\beta 2}\phi_{1}^{\pm}+c_{\beta 2}s_{\beta 1}\phi_{2}^{\pm}+s_{\beta 2}\phi_{3}^{\pm},\\ H_{2}^{\pm}&=(-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2})\phi_{1}^{\pm}+(c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2})\phi_{2}^{\pm}+(-c_{\beta 2}s_{\gamma 2})\phi_{3}^{\pm},\,\textrm{and}\\ H_{3}^{\pm}&=(-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2})\phi_{1}^{\pm}+(-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2})\phi_{2}^{\pm}+(c_{\beta 2}c_{\gamma 2})\phi_{3}^{\pm}.\end{split}$ (7)
Here, $\tan\beta_{1}=v_{2}/v_{1}$ and $\tan\beta_{2}=v_{3}/\sqrt{v_{1}^{2}+v_{2}^{2}}$. Finally, writing the mass terms for the CP-Odd Higgs in a similar fashion
$V_{n}^{mass}\supset\begin{pmatrix}n_{1}&n_{2}&n_{3}\end{pmatrix}\frac{\mathcal{M}^{2}_{n}}{2}\begin{pmatrix}n_{1}\\ n_{2}\\ n_{3}\end{pmatrix},$
We can diagonalize this matrix in a manner similar to the charged Higgs sector. The mass eigenstates are given by
$\begin{split}G_{0}&=(c_{\beta 1}c_{\beta 2})n_{1}+(c_{\beta 2}s_{\beta 1})n_{2}+(s_{\beta 2})n_{3},\\ A_{1}&=(-c_{\gamma 1}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 1})n_{1}+(c_{\beta 1}c_{\gamma 1}+s_{\beta 1}s_{\beta 2}s_{\gamma 1})n_{2}+(-c_{\beta 2}s_{\gamma 1})n_{3},\,\textrm{and}\\ A_{2}&=(-c_{\beta 1}c_{\gamma 1}s_{\beta 2}-s_{\beta 1}s_{\gamma 1})n_{1}+(-c_{\gamma 1}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 1})n_{2}+(c_{\beta 2}c_{\gamma 1})n_{3}.\end{split}$ (8)
The remaining $\lambda_{7,8,...12}$ terms of the potential are traded off in terms of the charged Higgs and CP-Odd Higgs masses,vevs and the mixing angles.
II.2 Yukawa Sector
Experimental limits place stringent constraints on the presence of tree-level Flavor Changing Neutral Currents (FCNCs), posing a significant challenge for multi-Higgs models. To address this, we impose Natural Flavor Conservation (NFC), a condition that restricts each Higgs doublet to couple with only one class of fermions. In the framework we adopt for the present work, we employ the Type-Z or democratic Yukawa structure, wherein the masses of the leptons, the down-type quarks, and the up-type quarks are generated by the $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ doublets respectively. This assignment not only satisfies the NFC criterion but also provides a natural mechanism to account for the observed hierarchy in fermion masses. We can thus write the Yukawa Lagrangian as
$\mathcal{L}_{Yukawa}=-[\bar{L}_{L}\Phi_{1}\mathcal{G}_{l}l_{R}+\bar{Q}_{L}\Phi_{2}\mathcal{G}_{d}d_{R}+\bar{Q}_{L}\tilde{\Phi}_{3}\mathcal{G}_{u}u_{R}+h.c],$ (9)
where the $\mathcal{G}_{f}$ are the Yukawa matrices. In terms of the fermion mass matrices they can be written as
$\mathcal{G}_{f}=\frac{\sqrt{2}\mathcal{M}_{f}}{v_{i}}.$
We work with a $Z_{3}$ symmetric potential as given in Eqn. 2 - for the Yukawa Lagrangian to remain invariant under the same, the right handed fermion fields transform as
$d_{R}\rightarrow\omega d_{R},\;\;\;\;l_{R}\rightarrow\omega^{2}l_{R},\;\;\;\;u_{R}\rightarrow u_{R}.$ (10)
II.3 Alignment-Limit
The discovery of a SM like Higgs boson necessitates that one of the CP-even scalars in the 3HDM be identified with it. This requirement imposes specific constraints on the parameters of the scalar potential and the mixing angles. In this work, we adopt the Regular Hierarchy configuration [8], wherein the lightest CP-even Higgs boson, $H_{1}$ is identified as the observed SM Higgs. The coupling of the $H_{1}$ with the SM gauge bosons in the democratic 3HDM is given by
$g_{HZZ}=\frac{ve^{2}(c_{\beta 2}c_{\alpha 2}\cos(\alpha_{1}-\beta_{1})+s_{\beta 2}s_{\alpha 2})}{2c_{w}^{2}s_{w}^{2}},$ (11)
and thus the alignment limit condition for the regular hierarchy reads
$c_{\beta_{2}}c_{\alpha_{2}}\cos(\alpha_{1}-\beta_{1})+s_{\beta_{2}}s_{\alpha_{2}}=1.$ (12)
Letting $k=\cos(\alpha_{1}-\beta_{1})$, we have
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$ and $\alpha_{2}=\beta_{2}+2n\pi$.
The case with $k=-1$ is equivalent to the current scenario up to an inversion, whereas the case with $k\neq 1$ represents a special configuration in which the entire vacuum expectation value (vev) is confined to the third doublet. Therefore, we proceed with our analysis by choosing $k=1$ [8].
II.4 Constraints
Herein, we list the various theoretical and experimental constraints that are directly relevant to understand the available parameter space of the model. For a more detailed study of how these constraints were imposed in this model, the reader is invited to consult Ref. [8].
  1. (i)
    Vacuum Stability: To ensure stability, the scalar potential must be bounded from below in all directions of the field space. This translates into a set of bounds on the scalar coupling parameters $\lambda_{i}$ which are listed below,
    $\begin{gathered}\lambda_{1}\geq 0,\quad\lambda_{2}\geq 0,\quad\lambda_{3}\geq 0,\\ \lambda_{4}+2\sqrt{\lambda_{1}\lambda_{2}}\geq 0,\quad\lambda_{5}+2\sqrt{\lambda_{1}\lambda_{3}}\geq 0,\quad\lambda_{6}+2\sqrt{\lambda_{2}\lambda_{3}}\geq 0,\\ \lambda_{4}+\lambda_{7}+2\sqrt{\lambda_{1}\lambda_{2}}\geq 0,\quad\lambda_{5}+\lambda_{8}+2\sqrt{\lambda_{1}\lambda_{3}}\geq 0,\quad\lambda_{6}+\lambda_{9}+2\sqrt{\lambda_{2}\lambda_{3}}\geq 0.\end{gathered}$ (13)
  2. (ii)
    Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies. Using the results of [17, 18] for unitarity Constraints in the $Z_{3}$ symmetric 3HDM, the condition to be imposed on the $21$ eigenvalues $\Lambda_{i}$ of the relevant scattering matrices is
    $|\Lambda_{i}|\leq 8\pi.$
  3. (iii)
    Perturbativity: That the model remains perturbatively calculable as a field theory is ensured by imposing on all the scalar couplings the condition $\lambda_{i}\leq|4\pi|.$
  4. (iv)
    EW Precision Tests: The Peskin-Takeuchi parameters $S$, $T$, and $U$ which capture the effects of new physics on gauge boson propagators impose stringent constraints on all extensions of the SM. We require that these parameters lie within the following allowed ranges:
    $S$ $=-0.02\pm 0.10,$
    $T$ $=0.03\pm 0.12,\,\textrm{and}$
    $U$ $=0.01\pm 0.11.$
  5. (v)
    BSM Higgs Boson Exclusion: The exclusion limits on the BSM scalar sector from direct searches conducted at the LHC, LEP, and Tevatron were computed at the $95\%$ Confidence Level (C.L.) using the HiggsBounds-6 framework, interfaced through the HiggsTools package [19] and imposed on the parameter space of the model.
  6. (vi)
    SM-Like Higgs Boson Discovery: The compatibility of the Higgs boson in the model with the discovered SM-like Higgs boson was ascertained using a goodness-of-fit test by HiggsSignals-5 via the HiggsTools package.
  7. (vii)
    Flavour Physics: The most stringent bound on the Branching Ratio ($\mathcal{BR}$) of the $B\rightarrow X_{s}\gamma$ decay using Next-to-Leading Order (NLO) calculations in QCD as discussed in [20, 21, 18] was used to constrain the model.
In what will follow in the subsequent sections, all the benchmark points chosen are those that are allowed after the imposition of these listed theoretical and experimental constraints. We now turn to the question of discovering the heavy Higgs bosons at the ILC.
III Phenomenological Analysis
To begin the study, we first list all possible modes of production and decay of the heavy scalars in this model at the ILC. We then perform a cut-based analysis for each of the considered production modes. Each channel is studied for two benchmark points ($BP$), with most cases analyzed in the fully hadronic final state.
III.1 Relevant Production Modes
We begin by listing all relevant production channels that can potentially uncover signals of a heavy Higgs in $e^{+}e^{-}$ collisions in Table 1. For each production mode, we have considered two sets of benchmark mass points, denoted as $BP\text{-}X$ and $BP\text{-}Y$, where $X$ and $Y$ are numerical labels assigned to clearly distinguish between different benchmark points across all production modes. These benchmarks satisfy all applied constraints and are thus suitable for a detailed analysis. Depending on the production mode, we consider either fully hadronic decays (FHD) or semi-leptonic decays (SLD) - the corresponding cross sections, branching ratios for the relevant signal channels, and the associated background processes will be discussed in the upcoming sections dedicated to the detailed phenomenological analysis.
Vertex Production Mode Final State Signal Benchmark
$SS$ $H_{2}A_{2}$ (III.2) $4b$ $BP\text{-}1$ , $BP\text{-}2$ : Table 2 $(FHD)$
$SSS$ $h_{1}H_{2}A_{2}$ (III.3) $6b$ $BP\text{-}3$ , $BP\text{-}4$ : Table 5 $(FHD)$
$SSV$ $H_{2}H_{2}Z$ (III.4) $4b+2j$ $BP\text{-}5$ , $BP\text{-}6$ : Table 9 $(FHD)$
$A_{2}A_{2}Z$ (III.5) $4b+2j$ $BP\text{-}7$ , $BP\text{-}8$ : Table 11 $(FHD)$
$H_{2}H_{2}Z$ (III.6) $4b+2\ell$ $BP\text{-}9$ , $BP\text{-}10$ : Table 14 $(SLD)$
$A_{2}A_{2}Z$ (III.7) $4b+2\ell$ $BP\text{-}11$ , $BP\text{-}12$ : Table 16 $(SLD)$
$H_{2}H_{2}^{\pm}W^{\mp}$ (III.8) $4b+4j$ $BP\text{-}13$ , $BP\text{-}14$ : Table 19 $(FHD)$
$A_{2}H_{2}^{\pm}W^{\mp}$ (III.9) $4b+4j$ $BP\text{-}15$ , $BP\text{-}16$ : Table 21 $(FHD)$
Table 1: Benchmark scenarios for different vertices and production modes, outlined with shorthand notations and corresponding table references.
The 3HDM, as mentioned in Sec. II, has multiple neutral and charged Higgs bosons. The choice of the specific channels in the table above is dictated by the parameter space that is allowed after all constraints have been imposed, and also demanding that the relevant couplings - and the associated cross-sections and branching ratios - be large enough that one can in principle probe them at colliders. This also explains why we have chosen a varied set of benchmark points.
III.2 $H_{2}A_{2}$ Production: $FHD$ Mode
For the $H_{2}A_{2}$ production mode, we consider the fully hadronic decay channel which in this case means that both $H_{2}$ and $A_{2}$ decay to $b\bar{b}$ resulting in a $4b$ final state. As previously mentioned, we study the efficacy of this channel (like all channels to follow) with two sets benchmark points ($BP$), with the corresponding branching ratios ($BR$) and total cross sections $(\sigma)$ provided in Table 2. In Fig. 1, we display a few illustrative Feynman diagrams for the production process. For the purpose of phenomenological analysis, we also identify the most relevant background processes—both reducible and irreducible—which are listed in Table 3.
[…]
Figure 1: A few illustrative Feynman diagrams for the $H_{2}A_{2}$ process.
$BP-1$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 340.46 409.61 274.85 469.73 376.15 400.29 $\sigma||BR$ 0.00598 0.75197 0.70063 0.0031528
$BP-2$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 287.95 457.92 209.50 425.93 330.31 441.07 $\sigma||BR$ 0.00782 0.74378 0.64988 0.0037801
Table 2: The upper portion of the table corresponds to the $BP\text{-}1$ mass points, $BR$ and total cross section ($\sigma$), while the lower portion corresponds to $BP\text{-}2$, for the $H_{2}A_{2}$ production mode and its fully hadronic decay.
Background Process Final State Cross Section($\sigma$) (pb)
$e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}$ $4b$ 0.006746
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ 0.06767
$e^{+}e^{-}\rightarrow t\bar{t}h\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ 0.0006778
Table 3: Relevant Standard Model background processes listed for $e^{+}e^{-}\to H_{2}A_{2}$, with $H_{2}\to b\bar{b}$ and $A_{2}\to b\bar{b}$.
[…]
Figure 2: The plots in the upper panel represent the selection cuts applied to $BP$-1, while those in the lower panel correspond to $BP$-2. Both sets of plots show the selection cuts on $p_{T}(b_{1})$ and $p_{T}(b_{2})$. The shaded regions indicate the selected ranges that are implemented in the cutflow chart.
The total cross sections for the two chosen benchmark points are of the same order, with only a small difference, making it reasonable to anticipate discovery under similar luminosity conditions. Furthermore, upon examining the observables after applying the $N(b)=4$ cut, the nature of their distinguishability remains comparable. Consequently, we have imposed identical cuts on the same observables, $p_{T}(b_{1})$ and $p_{T}(b_{2})$. The corresponding cutflow chart is presented in Table 4, where the significance is reported in the “S vs B” column using the formula $\frac{S}{\sqrt{S+B}}$, [^2] Throughout the discussion, the significance shown in the cutflow charts is consistently calculated using the formula $\frac{S}{\sqrt{S+B}}$. where $S$ and $B$ denote the number of signal and background events respectively. It is found that, at a luminosity of $100\,\mathrm{fb}^{-1}$, this channel exhibits a discovery potential of $5\sigma$ based on the implemented selection cuts. However, given that the kinematic properties of the four $b$ quarks are rather similar, it would not be possible to identify which pair belongs to the $H_{2}$ decay and which ones to the $A_{2}$, and thus, we do not attempt mass reconstruction. However, even identifying the $4b$ channel and imposing a minimal set of cuts helps us isolate BSM effects quite clearly in this scenario.
$BP-1$ : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 315 7509
$N(b)=4$ 58.82 85.11 4.903
SEL: $P_{T}(b_{1})>200$ 53.14 48.24 5.278
SEL: $P_{T}(b_{2})>150$ 47.27 35.20 5.206
$BP-2$ : Cuts Signal (S) Background (B) S vs B
Initial (no cut) 378 7509
$N(b)=4$ 64.79 85.11 5.292
SEL: $P_{T}(b_{1})>200$ 59.72 48.24 5.747
SEL: $P_{T}(b_{2})>150$ 54.81 35.2 5.779
Table 4: The upper table represents the cutflow chart for $BP-1$, while the lower table corresponds to $BP-2$. Both cases are considered at an integrated luminosity of $100\,\mathrm{fb}^{-1}$ in the analysis. In this setup, a discovery potential of $5\sigma$ is anticipated for the signal.
III.3 $h_{1}H_{2}A_{2}$ Production : $FHD$ Mode
This particular channel involving the production of $h_{1}H_{2}A_{2}$ and its subsequent decay into a fully hadronic mode with a $6b$ quark final state offers another promising avenue for exploration. The benchmark points considered here are referred to as $BP-3$ and $BP-4$ the corresponding masses, along with information regarding the branching ratios and total cross sections, are provided in Table 5.
A distinctive feature of this channel is its rare final state consisting of six $b$-quarks. The full production and decay chain is given below:
$e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}\quad(h_{1}\rightarrow b\bar{b})\quad(H_{2}\rightarrow b\bar{b})\quad(A_{2}\rightarrow b\bar{b})$
Due to the unusual final state compared to SM backgrounds, the signal may appear particularly clean, potentially reducing the need for cuts on observable distributions such as the momenta of b-quarks, especially if a strong pre-selection cut has already been applied. Based on the final state feature, a strong requirement of $N(b)=6$ is imposed [^3] We have implemented stringent preselection cuts here, where the multiplicity requirement exactly matches the partonic-level expectation. This does not significantly affect the signal efficiency, especially when weighed against the resulting reduction in background efficiency; otherwise, a slightly looser cut such as $N(b)\geq 6$ would have been considered. to effectively isolate signal events. In Table 6, we list both irreducible and reducible backgrounds that are relevant for this channel along with their cross-sections.
$BP-3$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $H_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 255.20 543.43 228.65 366.61 347.99 452.39 $\sigma||BR$ 0.0012179 0.54458 0.98325 0.981178 0.00063986
$BP-4$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $h_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
Mass (GeV) 125 269.54 470.15 215.39 522.60 311.28 451.15 $\sigma||BR$ 0.001619 0.54458 0.84123 0.82652 0.000613
Table 5: The upper portion of the table represents the mass points, $BR$ and total cross section corresponding to $BP-3$, whereas the lower portion corresponds to $BP-4$.
Background Process Final State Cross Section($\sigma$) (pb)
$e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}b\bar{b}$ $6b$ $5.886\times 10^{-6}$
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $0.06767$
$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $0.02346$
$e^{+}e^{-}\rightarrow t\bar{t}h\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ $6.778\times 10^{-4}$
$e^{+}e^{-}\rightarrow t\bar{t}Z\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ $2.774\times 10^{-4}$
Table 6: The relevant reducible and irreducible SM backgrounds for the 6 $b$ final state considered in this channel.
Given the highly distinctive nature of the final state, which is significantly different from typical SM backgrounds, it is reasonable to focus solely on counting the number of signal and background events based on a multiplicity cut of $N(b)=6$. At this stage, with the implementation of strong preselection cuts, it is highly unlikely that any of the SM backgrounds survive, as such a final state is relatively rare in $e^{+}e^{-}$ machines. The goal is to determine the range of luminosity required to effectively probe the production mechanism. The total cross sections for $BP-3$ and $BP-4$ are of similar magnitude, indicating that the required luminosity for discovery should also lie within a comparable range. We find that, for $BP-3$, an integrated luminosity of $1000\,\mathrm{fb}^{-1}$ and for $BP-4$, $1200\,\mathrm{fb}^{-1}$ are sufficient to achieve a discovery significance of $5\sigma$, based purely on the event count after applying the $N(b)=6$ selection. The corresponding selection cut and its effect on the event yield are summarized in the cutflow chart presented in Table 7.
$BP-3$:Cuts Signal (S) Background (B) S vs B
Initial (no cut) 639 92091
SEL: $N(b)=6$ 30.71 5.0 5.14
$BP-4$:Cuts Signal (S) Background (B) S vs B
Initial (no cut) 735 110509
SEL: $N(b)=6$ 32.28 6.0 5.217
Table 7: The upper portion of the table corresponds to the cutflow chart for $BP-3$ at $\mathcal{L}=1000\,fb^{-1}$, while the lower portion corresponds to $BP-4$ at $\mathcal{L}=1200\,fb^{-1}$.
III.4 $H_{2}H_{2}Z$ Production : $FHD$ Mode
We now extend our discussion to production modes that yield final states involving not only $b$-quarks but also additional light jets. In analyzing such final states - comprising various combinations of $b$-jets and light jets - it is essential to account for realistic detector-level effects. Although the parton-level topology predicts a specific number of jets, additional jets may be produced due to QCD radiation processes [22] such as initial state radiation (ISR), final state radiation (FSR), and jet fragmentation. Conversely, some jets may be too soft to be reconstructed or may fall outside the detector acceptance. Additionally, jets originating from hadronic decays of $W$ or $Z$ bosons may merge due to their collimated nature. To ensure robust signal retention while accounting for these effects, we adopt an inclusive jet multiplicity requirement of $N(j)\geq N_{\text{signal}}$, where $N_{\text{signal}}$ corresponds to the number of jets expected from the parton level signal topology. This approach captures events with additional QCD activity while maintaining consistency with the underlying signal hypothesis [^4] In previous scenarios such as $H_{2}A_{2}$ and $H_{1}H_{2}A_{2}$ production, we implemented cuts that matching with the exact parton level expectations. However, in the present case, doing so leads to a significant loss in signal efficiency relative to background rejection. To mitigate this, we adopt a more inclusive strategy that accounts for ISR, FSR, and jet fragmentation effects..
We now proceed with the analysis of the $H_{2}H_{2}Z$ production channel in the FHD mode, which results in a final state topology of $4b+2j$. The specific signal process is
$e^{+}e^{-}\to H_{2}H_{2}Z,\quad(H_{2}\to b\bar{b}),\quad(H_{2}\to b\bar{b}),\quad(Z\to jj).$
Although the signal contains exactly four $b$-quarks at the parton level, detector level reconstruction is impacted by $b$-tagging inefficiencies and possible contamination from gluon splitting processes. Accordingly, a requirement of $N(b)\geq 4$ is imposed to retain signal efficiency while suppressing background contributions, and is adopted along with a cut of $N(j)\geq 2$. Following the structure of the previously discussed channels, we consider two benchmark points for the analysis, denoted as $BP\text{-}5$ and $BP\text{-}6$, which are consistent with current experimental constraints. The relevant SM backgrounds for this process are listed in Table 8.555This same background table is also referenced in the $A_{2}A_{2}Z$ production case, as both processes in the FHD mode yield the same final state of $4b+2j$. Detailed information on the chosen benchmark points—including their masses, branching ratios, and total production cross sections—is presented in Table 9.
Background Process Final State Cross Section($\sigma$) (pb)
$e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}jj$ $4b+2j$ 0.00141
$e^{+}e^{-}\rightarrow W^{+}W^{-}b\bar{b}b\bar{b}\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ 0.0004943
$e^{+}e^{-}\rightarrow t\bar{t}h\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ 0.0006778
$e^{+}e^{-}\rightarrow t\bar{t}Z\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ 0.0002774
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ 0.06767
$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ 0.02346
$e^{+}e^{-}\rightarrow t\bar{t}jj\rightarrow b\bar{b}jjjjjj$ $2b+6j$ 0.005821
$e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ 0.001098
Table 8: The relevant SM backgrounds for the fully hadronic decay mode of $H_{2}H_{2}Z$ (Section III.4) and $A_{2}A_{2}Z$ (Section III.5) production.
BP-5:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma||BR$ 0.007175 0.815618 0.815618 0.548 0.002616
BP-6:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.548 0.002677
Table 9: The upper table corresponds to $BP-5$, and the lower table to $BP-6$, summarizing the benchmark mass points along with the associated $BR$ and total cross sections.
The selection cuts $N(b)\geq 4$ and $N(j)\geq 2$ are employed to preferentially isolate events consistent with the signal topology, thereby suppressing contributions from SM backgrounds. This approach, while simple in implementation, proves effective in enhancing the signal purity of the sample. Nevertheless, an important consequence of these inclusive number-based cuts is the significant overlap observed in the distributions of key kinematic observables between signal and background as can be evidenced from Fig. 3. As a result, no single observable emerges as a strong discriminator suitable for additional shape-based selection. Despite this limitation, the overall event yield remains sufficiently high. At an integrated luminosity of $300\,\mathrm{fb}^{-1}$, the signal reaches the threshold for discovery with a statistical significance of $5\sigma$. The cut-based selection criteria data is provided in Table 10.
Figure 3: Some sample kinematic distributions for the $H_{2}H_{2}Z$ decaying to a fully hadronic final state - as can be seen, the overlap of the signal and background distributions for much of the kinematic range does not permit a useful cut-based strategy in this case.
$BP-5$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1000 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 95.92 247.0 5.180
$BP-6$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1024 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 112.9 247.0 5.951
Table 10: The upper panel presents the cutflow summary for the BP-5 signal, while the lower panel shows the corresponding results for the BP-6 signal. Both cases assume an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
III.5 $A_{2}A_{2}Z$ Production : $FHD$ Mode
The $A_{2}A_{2}Z$ production channel and its subsequent decay topology are identical to the $H_{2}H_{2}Z$ case discussed in Section III.4. The corresponding SM backgrounds are listed in Table 8. The signal benchmarks, labeled as $BP\text{-}7$ and $BP\text{-}8$, correspond to different mass configurations. Detailed information on their mass values, BR’s and total production cross sections is provided in Table 11. The signal process is given by:
$e^{+}e^{-}\to A_{2}A_{2}Z,\quad(A_{2}\to b\bar{b}),\quad(A_{2}\to b\bar{b}),\quad(Z\to jj)$
A similar strategy is adopted for signal isolation in this channel, guided by the final state topology of $4b+2j$. In particular, number-based selection criteria of $N(b)\geq 4$ and $N(j)\geq 2$ are applied, consistent with the methodology outlined in the previous section. The cutflow summary for benchmark points $BP\text{-}7$ and $BP\text{-}8$ is provided in Table 12. Notably, even with only these inclusive number cuts, a favorable balance between signal efficiency and background suppression is achieved, highlighting the effectiveness of this minimal selection strategy.
BP-7:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.548 0.002341
BP-8:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.548 0.002709
Table 11: The upper table represents the mass points, BR and total cross section information for $BP-7$, whereas the lower table represents the same for $BP-8$.
$BP-7$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 895 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 108.43 247.0 5.752
$BP-8$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1036 30272
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 123.2 247.0 6.402
Table 12: The upper panel presents the cutflow summary for the BP-7 signal, while the lower panel shows the corresponding results for the BP-8 signal. Both cases are analyzed at an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
III.6 $H_{2}H_{2}Z$ Production : $SLD$ Mode
The $H_{2}H_{2}Z$ production channel discussed earlier in Section III.4 can also be considered with an alternative decay mode, wherein the $Z$ boson decays leptonically. The signal process is thus given by:
$e^{+}e^{-}\to H_{2}H_{2}Z,\quad(H_{2}\to b\bar{b}),\quad(H_{2}\to b\bar{b}),\quad(Z\to\ell^{+}\ell^{-})$
This leads to a distinct final state of $4b+2\ell$. One of the consequences could be a drop in the total cross section, as the $Z$ decaying hadronically yields a higher branching ratio, whereas its leptonic decay significantly reduces the total cross section due to a much lower branching ratio. Nevertheless, it must be compared against distinct and appropriate SM backgrounds. To proceed with the analysis, it is necessary to identify Standard Model backgrounds that can mimic this signature. The relevant backgrounds are listed in Table 13666The background processes listed apply equally to both $H_{2}H_{2}Z$ and $A_{2}A_{2}Z$ production channels in the semi-leptonic decay mode, as they yield identical final states.. Details of the two benchmark points, $BP\text{-}9$ and $BP\text{-}10$, including their masses, branching ratios, and cross sections, are provided in Table 14.
Background Process Final State Cross Section($\sigma$) (pb)
$e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}l^{+}l^{-}$ $4b+2l$ $1.893\times 10^{-5}$
[…]
Table 13: The relevant SM backgrounds for the semi-leptonic decay mode of $H_{2}H_{2}Z$ (Section III.6) and $A_{2}A_{2}Z$ (Section III.7) channels. The demand of at least two leptons necessitates the presence of SM gauge bosons that might also decay to neutrinos, and thus many of these backgrounds have sizable missing energy.
BP-9:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma||BR$ 0.007175 0.815618 0.815618 0.0673 0.0003212
BP-10:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.0673 0.0003288
Table 14: The upper portion of the table represents the mass points, BR, and total cross-section information for $BP-9$, while the lower portion represents the same for $BP-10$.
The analysis begins by applying number-based selection criteria, requiring $N(b)\geq 4$ and $N(\ell)\geq 2$, to facilitate effective reconstruction of the signal. Here, in this case as well, once the required number cuts are implemented, the remaining observable distributions for the signal and backgrounds become highly overlapping, making it difficult to distinguish between them. Also given that the BR for leptonic decays is significantly lower than that of the hadronic mode discussed earlier, a higher integrated luminosity is generally required to achieve comparable signal sensitivity. The cutflow summary for benchmark points $BP\text{-}9$ and $BP\text{-}10$ is presented in Table 15.
$BP-9$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 954 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 40.85 17.04 5.368
$BP-10$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 976 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 49.23 17.04 6.05
Table 15: The upper panel presents the cutflow summary for the BP-9 signal, while the lower panel shows the corresponding results for the BP-10 signal. Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
III.7 $A_{2}A_{2}Z$ Production : $SLD$ Mode
The $A_{2}A_{2}Z$ production channel followed by a semi-leptonic decay mode is analogous to the process discussed in Sec. III.6, yielding the same final state. Consequently, the relevant backgrounds remain unchanged and are listed in Table 13. Two benchmark mass points, denoted as $BP\text{-}11$ and $BP\text{-}12$, are considered, with their detailed properties provided in Table 16. The signal process is given by:
$e^{+}e^{-}\to A_{2}A_{2}Z,\quad(A_{2}\to b\bar{b}),\quad(A_{2}\to b\bar{b}),\quad(Z\to\ell^{+}\ell^{-})$
This process yields a final state of $4b+2\ell$. To enable effective signal reconstruction while suppressing Standard Model backgrounds, we impose number-based selection cuts of $N(b)\geq 4$ and $N(\ell)\geq 2$. The resulting cutflow summary, following the implementation of these selection criteria, is presented in Table 17.
BP-11:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.1 0.0002875
BP-12:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.0673 0.0003327
Table 16: The upper portion of the table represents the mass points, BR, and total cross-section values for $BP-11$, while the lower portion represents the same for $BP-12$.
$BP-11$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 853 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 55.04 17.04 6.482
$BP-12$: Cuts Signal (S) Background (B) S vs B
Initial (no cut) 988 22147
SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 67.28 17.04 7.326
Table 17: The upper panel presents the cutflow summary for the BP-11 signal, while the lower panel shows the corresponding results for the BP-12 signal. Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
With these findings, we now move on to discussing more non-trivial production modes and their prospects, particularly in probing vertices involving the presence of charged Higgs bosons, as we continue our phenomenological discussions.
III.8 $H_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
The $H_{2}H_{2}^{\pm}W^{\mp}$ production mode introduces a charged Higgs boson in the production mode [^7] This scenario bears a close resemblance to the $A_{2}H_{2}^{\pm}W^{\mp}$ channel in the FHD mode discussed in Subsection III.9.. This particular channel features a more non-trivial and complex final state, with a higher jet multiplicity than those discussed so far. The processe culminate to a final state topology of $4b+4j$. While this signature is relatively rare in SM processes, certain backgrounds can still mimic it due to effects such as mistagging and QCD jet production. We identify and compile a unified set of relevant SM backgrounds, which are listed in Table 18.888This background table is also applicable to the subsequent analysis of the $A_{2}H_{2}^{\pm}W^{\mp}$ process in the FHD mode. The signal process for the $H_{2}H_{2}^{\pm}W^{\mp}$ production channel proceeds through the following decay chain:
$e^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp},\quad(H_{2}\to b\bar{b}),\quad(H_{2}^{\pm}\to t\bar{b}/\bar{t}b\to W^{+}b\bar{b}/W^{-}b\bar{b}\to jjb\bar{b}),\quad(W^{\mp}\to jj)$
Two representative benchmark points, $BP\text{-}13$ and $BP\text{-}14$, are considered in this study. Their mass configurations, BR and total production cross sections are summarized in Table 19. Given the complexity of the $4b+4j$ final state and the presence of fully hadronic decays, our selection strategy aims to keep as much signal as possible while reducing the background as much as we can. To this end, we apply the number-based selection cuts $N(b)\geq 4$ and $N(j)\geq 4$. These cuts are well-aligned with the parton-level expectations of the signal process and are effective in suppressing SM backgrounds, particularly those lacking sufficient $b$-quark content. Although the hadronic nature of the decay and the low signal cross section present challenges, we explore the discovery potential of this channel under a high-luminosity scenario. Our analysis indicates that, with an integrated luminosity of $2000~{}\text{fb}^{-1}$, the signal can be clearly discovered with a significance of $5\sigma$. The cutflow chart corresponding to this analysis, based on cut-based selections, is shown in Table 20.
Background Process Final State Cross Section($\sigma$) (pb)
$e^{+}e^{-}\rightarrow ZZb\bar{b}b\bar{b}\rightarrow jjjjb\bar{b}b\bar{b}$ $4b+4j$ $2.089\times 10^{-7}$
$e^{+}e^{-}\rightarrow W^{+}W^{-}b\bar{b}b\bar{b}\rightarrow jjjjb\bar{b}b\bar{b}$ $4b+4j$ $4.943\times 10^{-4}$
$e^{+}e^{-}\rightarrow t\bar{t}h\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ $6.778\times 10^{-4}$
$e^{+}e^{-}\rightarrow t\bar{t}Z\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ $2.774\times 10^{-4}$
$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $0.06767$
$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $0.02346$
$e^{+}e^{-}\rightarrow t\bar{t}jj\rightarrow b\bar{b}jjjjjj$ $2b+6j$ $0.005821$
$e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ $0.001098$
Table 18: Relevant Standard Model backgrounds for the $H_{2}H_{2}^{\pm}W^{\mp}$ (Section III.8) and $A_{2}H_{2}^{\pm}W^{\mp}$ (Section III.9) production channels in the fully hadronic decay ($FHD$) mode, featuring a final state topology of $4b+4j$.
$BP-13$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 224.16 359.96 377.98 474.54 326.93 500.67 $\sigma||BR$ 0.01095 0.8232 0.1575 1.0 0.6 0.001022
$BP-14$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 245.16 505.60 358.58 463.76 335.25 395.24 $\sigma||BR$ 0.01035 0.9159 0.1889 1.0 0.6 0.00129
Table 19: The table provides information on $H_{2}H_{2}^{\pm}W^{\mp}$ production and its subsequent decay into a $4b+4j$ final state. The upper portion of the table corresponds to $BP-13$, while the lower portion corresponds to $BP-14$.
$BP-13:$Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1533 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 104.0 302.8 5.15
$BP-14:$Cuts Signal (S) Background (B) S vs B
Initial (no cut) 1934 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 136.4 302.8 6.51
Table 20: Cutflow charts for benchmark points $BP$-$13$ (upper table) and $BP$-$14$ (lower table). Both signals are analyzed at an integrated luminosity of $2000~{}\text{fb}^{-1}$.
III.9 $A_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
The $A_{2}H_{2}^{\pm}W^{\mp}$ production mode shares its final state topology with the previously discussed $H_{2}H_{2}^{\pm}W^{\mp}$ process, both culminating in a fully hadronic signature of $4b+4j$. This overlap in decay structure naturally suggests a unified treatment of Standard Model backgrounds, particularly those capable of mimicking this rare final state via mistagging. The common background processes relevant to both channels are listed in Table 18. The full production and decay chain under consideration is:
$e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp},\quad(A_{2}\to b\bar{b}),\quad(H_{2}^{\pm}\to t\bar{b}/\bar{t}b\to W^{+}b\bar{b}/W^{-}b\bar{b}\to jjb\bar{b}),\quad(W^{\mp}\to jj)$
We select two benchmark points—denoted as $BP\text{-}15$ and $BP\text{-}16$—that satisfy existing phenomenological constraints while offering representative signal kinematics. Their detailed properties, including masses, decay branching fractions, and production cross sections, are provided in Table 21. To extract the signal from the multijet-dominated environment, we impose a set of number-based selection cuts consistent with the parton-level topology: $N(b)\geq 4$ and $N(j)\geq 4$. This inclusive approach accounts for realistic detector effects, such as initial and final-state radiation and jet fragmentation, while maintaining strong discriminatory power against backgrounds lacking sufficient $b$-jet multiplicity. Despite the inherently suppressed cross sections and the challenging hadronic environment, we investigate whether this channel can yield meaningful statistical significance at a high-luminosity collider setup. Our results, summarized in the cutflow Table 22, indicate that with an integrated luminosity of $2000~{}\text{fb}^{-1}$, the signal can be probed with a discovery-level sensitivity of $5\sigma$.
BP-15:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 365.69 370.25 222.74 366.90 309.51 437.50 $\sigma||BR$ 0.01111 0.96079 0.28799 1.0 0.6 0.002214
BP-16:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
Mass (GeV) 125 286.48 386.58 204.09 360.96 300.06 401.01 $\sigma||BR$ 0.011191 0.95317 0.24065 1.0 0.6 0.001848
Table 21: The upper table presents the mass points, branching ratios ($BR$), and total cross section for $BP-15$, while the lower table provides the corresponding information for $BP-16$.
$BP-15:$ Cuts Signal (S) Background (B) S vs B
Initial (no cut) 3320 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 110.24 302.8 5.43
$BP-16:$ Cuts Signal (S) Background (B) S vs B
Initial (no cut) 2772 298496
SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 143.73 302.8 6.8
Table 22: The upper portion of the table presents the cutflow chart for $BP\text{-}15$, while the lower portion corresponds to $BP\text{-}16$. At an integrated luminosity of $2000~{}\text{fb}^{-1}$.
With all these channel surveys carried out, we now move on to the Discussion and Conclusion section section IV, where we highlight the relevant discussions and outline future outlooks that could be considered in experiments such as the ILC.
IV Discussion and Conclusions
Various present and future collider experiments offer distinct capabilities to probe new physics, each with their own unique strengths and experimental setups. The International Linear Collider, in particular, provides a clean experimental environment and energy tunability, which make it especially promising for precision studies of extended Higgs sectors [16, 23, 24]. Motivated by these advantages, we have investigated the potential of the ILC operating at a center-of-mass energy of $\sqrt{s}=1000$ GeV to probe physics beyond the SM through an extension of the scalar sector. In this study, we focus on the 3HDM, which provides scope for a rich phenomenology due to its enlarged scalar spectrum and expanded parameter space. The designed framework in the work respects theoretical consistency and complies with current experimental constraints. Our goal was to identify promising search strategies sensitive to interaction vertices characteristic of new physics, and to explore benchmark scenarios where viable production cross sections could be obtained, potentially leading to observable signals.
We performed a detailed analysis of several production channels at the ILC, including $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $e^{+}e^{-}\to A_{2}A_{2}Z$, $e^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp}$, $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp}$, and $e^{+}e^{-}\to H_{1}H_{2}A_{2}$, which yield final states containing multiple jets, leptons, and b-quarks, making them ideal channels to investigate the extended Higgs sector. For each channel, we selected representative benchmark points and conducted detailed phenomenological analysis. It was observed that, with the exception of the $e^{+}e^{-}\to H_{2}A_{2}$ channel, the application of additional kinematic variables such as $p_{T}(j_{1})$ or $p_{T}(b_{1})$ beyond the required object multiplicity cuts offers limited improvement in signal sensitivity. This is due to significant overlap between signal and background distributions after basic selection, making further discrimination through such observables less effective.
This study presents a phenomenological survey of several channels through which new physics interaction vertices can be probed. In all the processes considered, at least two new particles contribute to the final state. For example, in $e^{+}e^{-}\to H_{2}A_{2}$, both $H_{2}$ and $A_{2}$ decay into $b$-quark pairs with overlapping kinematics due to their similar masses, making it difficult to reconstruct them individually from $b$-jet pairings. Nonetheless, an excess of events can still reveal the presence of such interactions. Similar challenges arise in all other channels. While addressing these combinatorial issues is beyond the scope of this work, future studies may explore dedicated strategies for improving signal reconstruction in such scenarios [25, 26]. However, the channels considered in our study do exhibit promising discovery potential at various luminosity thresholds. With the application of suitable number-based selection cuts and transverse momentum requirements on the leading and subleading $b$-quarks, the $H_{2}A_{2}$ channel yields a $5\sigma$ discovery significance at an integrated luminosity of $100~{}\text{fb}^{-1}$ for both benchmark points BP-1 and BP-2. For the $h_{1}H_{2}A_{2}$ channel, discovery is achievable at $1000~{}\text{fb}^{-1}$ for BP-3 and at $1200~{}\text{fb}^{-1}$ for BP-4. Similarly, the $H_{2}H_{2}Z$/$A_{2}A_{2}Z$ channel in the FHD mode requires approximately $300~{}\text{fb}^{-1}$ for discovery, whereas in the SLD mode, it requires about $2000~{}\text{fb}^{-1}$. Finally, the channel involving a charged Higgs, $H_{2}H_{2}^{\pm}W^{\mp}$/$A_{2}H_{2}^{\pm}W^{\mp}$, in the FHD mode demands a luminosity of up to $2000~{}\text{fb}^{-1}$ to reach discovery sensitivity. These results demonstrate that several production channels hinting at new physics can achieve a $5\sigma$ discovery at this collider setup.
Overall, this study highlights promising directions for probing extended Higgs sectors at the ILC and providing the information of channel-wise analysis. We have compiled all the production channels considered in this work along with their respective decay modes, as summarized in Table 1. The discovery prospects of each production mode have been evaluated under different luminosity scenarios, as shown through detailed cutflow analyses [^9] Most of the cutflow steps involve object multiplicity requirements designed to match the parton-level expectations.. The resulting signal significances for various channels, compiled across varying luminosities, are presented as the main outcome of this phenomenological survey in Figure 4.
Figure 4: Signal significance vs luminosity for all considered channels.
V Acknowledgements
A.K. acknowledges the support from the Director’s Fellowship at IIT Gandhinagar. G.B.K. acknowledges the initial financial support received under the Prime Minister’s Research Fellowship (Grant No. PMRF-192002-1802), awarded by the Ministry of Education, Government of India.

### v2

@@
-3HDM at the ILC
+ILC Phenomenology of the $Z_{3}$ symmetric Type-Z Three Higgs Doublet Model
 Abstract
@@
 Applying a comprehensive set of theoretical and experimental constraints, we identify promising new physics signals with sufficiently large production cross-sections.
-Our analysis shows that $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $A_{2}A_{2}Z$, $H_{2}H_{2}^{\pm}W^{\mp}$, $A_{2}H_{2}^{\pm}W^{\mp}$ and $H_{1}H_{2}A_{2}$ are among the most sensitive channels to probe this extended Higgs sector.
+Our analysis shows that $^{+}e^{-}\to H_{2}A_{2}, e+e$$^{+}e^{-}\to H_{2}H_{2}Z$, $A_{2}A_{2}Z$, $H_{2}H_{2}^{\pm}W^{\mp}$, $A_{2}H_{2}^{\pm}W^{\mp}$ and $H_{1}H_{2}A_{2}$ are among the most sensitive channels to probe this extended Higgs sector.
 We demonstrate that a future ILC would offer a powerful platform to test these interactions and discover these heavier Higgs bosons thus providing evidence of physics beyond the Standard Model.
@@
 One of the many ways the theoretical framework of the SM can be extended is by adding additional scalar sectors, such as singlet [3, 4] and doublet extensions [5], and their generalization to multi-Higgs doublet models [6, 7].
-Of late, along these lines, the 3HDM has emerged as a potentially interesting and promising source of investigation [8, 9].
+Along these lines, the 3HDM [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] has also been a long-standing source of investigation.
 These models have been key in addressing unresolved questions, such as providing new sources of CP violation which are crucial for explaining baryogenesis [18], enabling strong first-order electroweak phase transitions, a necessary condition for successful electroweak baryogenesis [19].
 In certain realizations, the extended scalar sector can also offer viable dark matter candidates through neutral scalars [20].
-Beyond cosmological implications, such models predict a richer Higgs spectrum with distinct collider signatures, offering promising discovery opportunities at current and future colliders like the LHC and ILC.
+Beyond cosmological implications, such models predict a richer Higgs spectrum with distinct collider signatures, offering promising discovery opportunities at current and future colliders like the LHC and International Linear Collider (ILC) [21].
 While models with two Higgs doublets have been extensively studied [5], the 3HDM offers an even richer framework for addressing fundamental questions beyond the Standard Model.
 Furthermore, the additional scalar degrees of freedom enable more flexible Yukawa structures, providing potential explanations for the observed patterns of fermion masses and mixing [22].
-From a phenomenological perspective, the extended scalar spectrum enables a variety of novel collider signatures including multiple charged and neutral Higgs bosons that can be probed at future experiments such as the International Linear Collider (ILC) or the Future Circular Collider (FCC-ee).
-In this paper, we explore the phenomenological aspects of the 3HDM from the standpoint of the future ILC experiment.
+From a phenomenological perspective, the extended scalar spectrum enables a variety of novel collider signatures including multiple charged and neutral Higgs bosons that can be probed at future experiments such as the ILC or the Future Circular Collider (FCC-ee) [23].
+For a complete review of BSM scenarios with extended scalar sectors, the reader is invited to consult Ref.
+[13].
+In this paper, we explore the phenomenological aspects of the $Z_{3}$ symmetric Type Z 3HDM from the standpoint of the future ILC experiment.
 As the ILC global project proposal [24, 25] continues to evolve and gain traction within the community, it is timely and important to investigate the avenues through which various Beyond Standard Model (BSM) scenarios could manifest in its setup.
@@
 We focus on channels that are expected to have reasonable cross sections while operating with benchmark points allowed by all existing theoretical and experimental constraints.
-The paper is organized as follows: in Sec.
-II, we provide a brief overview of the 3HDM concentraing on the scalar structure, and list the theoretical and experimental constraints on the model.
-The phenomenological analysis of the various relevant channels at the ILC is presented in Sec.
-III, and we present our conclusions in Sec.
-IV.
+The paper is organized as follows: in section II, we provide a brief overview of the 3HDM concentraing on the scalar structure, and list the theoretical and experimental constraints on the model.
+The phenomenological analysis of the various relevant channels at the ILC is presented in section III, and we present our conclusions in section IV.
 II Model and Constraints
@@
 $\Phi_{k}=\begin{pmatrix}\phi_{k}^{+}\\ \frac{v_{k}+p_{k}+in_{k}}{\sqrt{2}}\end{pmatrix}$, (1)
-where $k=1,2,3$. We consider a $Z_{3}$ symmetric potential under which the Higgs fields transform as
+where $k=1,2,3$. We consider a $Z_{3}$ symmetric potential [27, 28, 29] under which the Higgs fields transform as
 $\Phi_{1}\rightarrow\omega\Phi_{1},\;\;\Phi_{2}\rightarrow\omega^{2}\Phi_{2},\;\textrm{and}\;\Phi_{3}\rightarrow\Phi_{3},$ (2)
-where $\omega=e^{2\pi i/3}$ are the cube roots of unity. The most general $SU(2)_{L}\times U(1)_{Y}$ invariant potential preserving the $Z_{3}$ symmetry is given by
-$\begin{split}V&=m_{11}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+m_{22}^{2}(\Phi_{2}^{\dagger}\Phi_{2})+m_{33}^{2}(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{1}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\lambda_{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}+\lambda_{3}(\Phi_{3}^{\dagger}\Phi_{3})^{2}\\ &+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\lambda_{5}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{3}^{\dagger}\Phi_{3})+\lambda_{6}(\Phi_{2}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{3})\\ &+\lambda_{7}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})+\lambda_{8}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{1})+\lambda_{9}(\Phi_{2}^{\dagger}\Phi_{3})(\Phi_{3}^{\dagger}\Phi_{2})\\ &+[\lambda_{10}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{1}^{\dagger}\Phi_{3})+\lambda_{11}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{3}^{\dagger}\Phi_{2})+\lambda_{12}(\Phi_{1}^{\dagger}\Phi_{3})(\Phi_{2}^{\dagger}\Phi_{3})+h.c.].\\ &\end{split}$ (3)
+where $\omega=e^{2\pi i/3}$ are the nontrivial cubic roots of unity. The most general $SU(2)_{L}\times U(1)_{Y}$ invariant potential preserving the $Z_{3}$ symmetry is given by
+[…]
 Here, in general $\lambda_{1,...,9}$ are real parameters, while $\lambda_{10},\lambda_{11},\lambda_{12}$ can be complex.
 However, here we disregard the complex phases of the potential to prevent mixing between the CP-odd and CP-even states [8].
-The relevant mass terms for the CP-even Higgs are deduced from the scalar potential in Eqn.
-3, and can be written as
+Consequently, the scalar potential is CP conserving, with no explicit CP violation.
+Further, in what follows, we also assume there is no spontaneous CP violation as well, i.e., we assume no relative phases in the vacuum expectation values are generated through minimization.
+The relevant mass terms for the CP-even Higgs are deduced from the scalar potential in Equation 3, and can be written as
 $V_{p}^{mass}\supset\begin{pmatrix}p_{1}&p_{2}&p_{3}\end{pmatrix}\frac{\mathcal{M}^{2}_{S}}{2}\begin{pmatrix}p_{1}\\ p_{2}\\ p_{3}\end{pmatrix}$,
 As a real symmetric matrix, it can be diagonalized through an orthogonal transformation characterized by the matrix $O_{\alpha}$. The diagonalization condition is given by
-$O_{\alpha}.\mathcal{M}^{2}_{S}.O_{\alpha}^{T}=\begin{pmatrix}m_{H1}^{2}&0&0\\ 0&m_{H2}^{2}&0\\ 0&0&m_{H3}^{2}\end{pmatrix}$ (4)
+$O_{\alpha}\mathcal{M}^{2}_{S}O_{\alpha}^{T}=\begin{pmatrix}m_{H1}^{2}&0&0\\ 0&m_{H2}^{2}&0\\ 0&0&m_{H3}^{2}\end{pmatrix}$ (4)
 The explicit relations of $\lambda_{1,2,..,6}$ can be expressed in terms of the masses of the three CP-even Higgs bosons and three mixing angles.
@@
 $\begin{split}H_{1}&=c_{\alpha_{2}}c_{\alpha_{1}}p_{1}+c_{\alpha_{2}}s_{\alpha_{1}}p_{2}+s_{\alpha_{2}}p_{3},\\ H_{2}&=-(c_{\alpha_{3}}s_{\alpha_{1}}+s_{\alpha_{3}}s_{\alpha_{2}}c_{\alpha_{1}})p_{1}+(c_{\alpha_{3}}c_{\alpha_{1}}-s_{\alpha_{3}}s_{\alpha_{2}}s_{\alpha_{1}})p_{2}+(s_{\alpha_{3}}c_{\alpha_{2}})p_{3},\,\textrm{and}\\ H_{3}&=(s_{\alpha_{3}}s_{\alpha_{1}}-c_{\alpha_{3}}s_{\alpha_{2}}c_{\alpha_{1}})p_{1}-(s_{\alpha_{3}}c_{\alpha_{1}}+c_{\alpha_{3}}s_{\alpha_{2}}s_{\alpha_{1}})p_{2}+(c_{\alpha_{3}}c_{\alpha_{2}})p_{3}.\end{split}$ (5)
-The mass terms for the charged Higgses can similarly be extracted from the scalar potential given in Eqn. 3:
+The mass terms for the charged Higgses can similarly be extracted from the scalar potential given in Equation 3:
 $V_{C}^{mass}\supset\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\mathcal{M}^{2}_{\phi^{\pm}}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix}$,
@@
 where
-$(O_{\gamma 2}.O_{\beta})^{\dagger}=\begin{pmatrix}c_{\beta 1}c_{\beta 2}&-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2}\\ c_{\beta 2}s_{\beta 1}&c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2}\\ s_{\beta 2}&-c_{\beta 2}s_{\gamma 2}&c_{\beta 2}c_{\gamma 2}\end{pmatrix}.$ (6)
+$(O_{\gamma 2}O_{\beta})^{\dagger}=\begin{pmatrix}c_{\beta 1}c_{\beta 2}&-c_{\gamma 2}s_{\beta 1}+c_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\beta 1}c_{\gamma 2}s_{\beta 2}-s_{\beta 1}s_{\gamma 2}\\ c_{\beta 2}s_{\beta 1}&c_{\beta 1}c_{\gamma 2}+s_{\beta 1}s_{\beta 2}s_{\gamma 2}&-c_{\gamma 2}s_{\beta 1}s_{\beta 2}+c_{\beta 1}s_{\gamma 2}\\ s_{\beta 2}&-c_{\beta 2}s_{\gamma 2}&c_{\beta 2}c_{\gamma 2}\end{pmatrix}$.
+(6)
 The gauge eigenstates can be represented in terms of mass eigenstates as
@@
 In the framework we adopt for the present work, we employ the Type-Z or democratic Yukawa structure, wherein the masses of the leptons, the down-type quarks, and the up-type quarks are generated by the $\Phi_{1}$, $\Phi_{2}$, and $\Phi_{3}$ doublets respectively.
-This assignment not only satisfies the NFC criterion but also provides a natural mechanism to account for the observed hierarchy in fermion masses.
 We can thus write the Yukawa Lagrangian as
@@
 $\mathcal{G}_{f}=\frac{\sqrt{2}\mathcal{M}_{f}}{v_{i}}.$
-We work with a $Z_{3}$ symmetric potential as given in Eqn. 2 - for the Yukawa Lagrangian to remain invariant under the same, the right handed fermion fields transform as
+We work with a $Z_{3}$ symmetric potential as given in Equation 2 - for the Yukawa Lagrangian to remain invariant under the same, the right handed fermion fields transform as
 $d_{R}\rightarrow\omega d_{R},\;\;\;\;l_{R}\rightarrow\omega^{2}l_{R},\;\;\;\;u_{R}\rightarrow u_{R}.$ (10)
@@
 Letting $k=\cos(\alpha_{1}-\beta_{1})$, we have
-  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$ and $\alpha_{2}=\beta_{2}+2n\pi$.
-The case with $k=-1$ is equivalent to the current scenario up to an inversion, whereas the case with $k\neq 1$ represents a special configuration in which the entire vacuum expectation value (vev) is confined to the third doublet.
+  • $k=1\implies\alpha_{1}=\beta_{1}$ and $\alpha_{2}=\beta_{2}$.
+The case with k=1k=-1 is equivalent to the current scenario up to an inversion, whereas the case with $k\neq 1$ represents a special configuration in which the entire vacuum expectation value (vev) is confined to the third doublet.
 Therefore, we proceed with our analysis by choosing $k=1$ [8].
@@
     Vacuum Stability: To ensure stability, the scalar potential must be bounded from below in all directions of the field space.
+[^1] We remark here that the complete necessary and sufficient conditions for the $Z_{3}$ symmetric 3HDM are not yet known.
+The conditions listed in Eq.
+13 should instead be understood as a set of sufficient stability conditions obtained by requiring positivity of the quartic potential along a representative set of field directions in the scalar field space.
+These conditions are analogous to the well-known copositivity-inspired constraints [30], used in multi-Higgs models when the full analytic BFB solution is unavailable.
+While they do not exhaust all possible directions in field space, they guarantee stability along all tested directions and therefore ensure that the scalar potential is bounded from below in a physically relevant region of parameter space.
+Also see Ref [28, 31].
 This translates into a set of bounds on the scalar coupling parameters $\lambda_{i}$ which are listed below,
@@
     Unitarity: The unitarity constraints are necessary to ensure that the theory remains predictive at high energies.
-Using the results of [17, 18] for unitarity Constraints in the $Z_{3}$ symmetric 3HDM, the condition to be imposed on the $21$ eigenvalues $\Lambda_{i}$ of the relevant scattering matrices is
+Using the results of [32, 33] for unitarity Constraints in the $Z_{3}$ symmetric 3HDM, the condition to be imposed on the 2121 eigenvalues $\Lambda_{i}$ of the relevant scattering matrices is
     $|\Lambda_{i}|\leq 8\pi.$
@@
 We begin by listing all relevant production channels that can potentially uncover signals of a heavy Higgs in $e^{+}e^{-}$ collisions in Table 1.
-For each production mode, we have considered two sets of benchmark mass points, denoted as $BP\text{-}X$ and $BP\text{-}Y$, where $X$ and $Y$ are numerical labels assigned to clearly distinguish between different benchmark points across all production modes.
+For each production mode, we have considered two sets of benchmark mass points, denoted as BPX and BPY, where X and Y are numerical labels assigned to clearly distinguish between different benchmark points across all production modes.
 These benchmarks satisfy all applied constraints and are thus suitable for a detailed analysis.
@@
 Vertex Production Mode Final State Signal Benchmark
-$SS$ $H_{2}A_{2}$ (III.2) $4b$ $BP\text{-}1$ , $BP\text{-}2$ : Table 2 $(FHD)$
-$SSS$ $h_{1}H_{2}A_{2}$ (III.3) $6b$ $BP\text{-}3$ , $BP\text{-}4$ : Table 5 $(FHD)$
-$SSV$ $H_{2}H_{2}Z$ (III.4) $4b+2j$ $BP\text{-}5$ , $BP\text{-}6$ : Table 9 $(FHD)$
-$A_{2}A_{2}Z$ (III.5) $4b+2j$ $BP\text{-}7$ , $BP\text{-}8$ : Table 11 $(FHD)$
-$H_{2}H_{2}Z$ (III.6) $4b+2\ell$ $BP\text{-}9$ , $BP\text{-}10$ : Table 14 $(SLD)$
-$A_{2}A_{2}Z$ (III.7) $4b+2\ell$ $BP\text{-}11$ , $BP\text{-}12$ : Table 16 $(SLD)$
-$H_{2}H_{2}^{\pm}W^{\mp}$ (III.8) $4b+4j$ $BP\text{-}13$ , $BP\text{-}14$ : Table 19 $(FHD)$
-$A_{2}H_{2}^{\pm}W^{\mp}$ (III.9) $4b+4j$ $BP\text{-}15$ , $BP\text{-}16$ : Table 21 $(FHD)$
+SSSS $H_{2}A_{2}$ (III.2) $4b$ BP1 : Table 2 $(FHD)$
+SSSSSS $h_{1}H_{2}A_{2}$ (III.3) $6b$ BP2 : Table 5 $(FHD)$
+SSVSSV $H_{2}H_{2}Z$ (III.4) $4b+2j$ BP3 : Table 9 $(FHD)$
+$A_{2}A_{2}Z$ (III.5) $4b+2j$ BP4 : Table 11 $(FHD)$
+$H_{2}H_{2}Z$ (III.6) $4b+2\ell$ BP5 : Table 14 $(SLD)$
+$A_{2}A_{2}Z$ (III.7) $4b+2\ell$ BP6 : Table 16 $(SLD)$
+$H_{2}H_{2}^{\pm}W^{\mp}$ (III.8) $4b+4j$ BP7 : Table 19 $(FHD)$
+$A_{2}H_{2}^{\pm}W^{\mp}$ (III.9) $4b+4j$ BP8 : Table 20 $(FHD)$
 Table 1: Benchmark scenarios for different vertices and production modes, outlined with shorthand notations and corresponding table references.
-The 3HDM, as mentioned in Sec.
-II, has multiple neutral and charged Higgs bosons.
+The 3HDM, as mentioned in section II, has multiple neutral and charged Higgs bosons.
 The choice of the specific channels in the table above is dictated by the parameter space that is allowed after all constraints have been imposed, and also demanding that the relevant couplings - and the associated cross-sections and branching ratios - be large enough that one can in principle probe them at colliders.
@@
 For the $H_{2}A_{2}$ production mode, we consider the fully hadronic decay channel which in this case means that both $H_{2}$ and $A_{2}$ decay to $b\bar{b}$ resulting in a $4b$ final state.
-As previously mentioned, we study the efficacy of this channel (like all channels to follow) with two sets benchmark points ($BP$), with the corresponding branching ratios ($BR$) and total cross sections $(\sigma)$ provided in Table 2.
-In Fig.
-1, we display a few illustrative Feynman diagrams for the production process.
+As previously mentioned, we study the efficacy of this channel (like all channels to follow) with a benchmark point (BP), with the corresponding branching ratios (BR) and total cross sections $(\sigma)$ provided in Table 2.
+In Figure 1, we display a few illustrative Feynman diagrams for the production process.
 For the purpose of phenomenological analysis, we also identify the most relevant background processes—both reducible and irreducible—which are listed in Table 3.
@@
 Figure 1: A few illustrative Feynman diagrams for the $H_{2}A_{2}$ process.
-$BP-1$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
-Mass (GeV) 125 340.46 409.61 274.85 469.73 376.15 400.29 $\sigma||BR$ 0.00598 0.75197 0.70063 0.0031528
-$BP-2$:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
-Mass (GeV) 125 287.95 457.92 209.50 425.93 330.31 441.07 $\sigma||BR$ 0.00782 0.74378 0.64988 0.0037801
-Table 2: The upper portion of the table corresponds to the $BP\text{-}1$ mass points, $BR$ and total cross section ($\sigma$), while the lower portion corresponds to $BP\text{-}2$, for the $H_{2}A_{2}$ production mode and its fully hadronic decay.
-Background Process Final State Cross Section($\sigma$) (pb)
+BP1:Particles $\rightarrow$ $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}A_{2}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
+Mass (GeV) 125 340.46 409.61 274.85 469.73 376.15 400.29 $\sigma\times BR$ 0.00598 0.75197 0.70063 0.0031528
+Table 2: The table corresponds to the BP1 mass points, BR and total cross section ($\sigma$), for the $H_{2}A_{2}$ production mode and its fully hadronic decay.
+Background Process Final State Cross Section ($\sigma$) [pb]
 $e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}$ $4b$ 0.006746
@@
 $e^{+}e^{-}\rightarrow t\bar{t}h\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ 0.0006778
-Table 3: Relevant Standard Model background processes listed for $e^{+}e^{-}\to H_{2}A_{2}$, with $H_{2}\to b\bar{b}$ and $A_{2}\to b\bar{b}$.
-[…]
-Figure 2: The plots in the upper panel represent the selection cuts applied to $BP$-1, while those in the lower panel correspond to $BP$-2.
+Table 3: Relevant Standard Model background processes for $e^{+}e^{-}\to H_{2}A_{2}$, with $H_{2}\to b\bar{b}$ and $A_{2}\to b\bar{b}$.
+To isolate the signal in our analysis, we impose a stringent pre-selection cut of N(b)=4N(b)=4222An alternative strategy could involve a more inclusive requirement such as $N(b)\geq 4$.
+In realistic detector environments, additional jets may arise due to initial-state radiation (ISR), final-state radiation (FSR), and jet fragmentation.
+Such effects can result in a higher jet multiplicity than the parton-level expectation.
+Therefore, using $N(b)\geq 4$ may be beneficial in scenarios where a strict cut significantly reduces signal efficiency.
+However, in the present case, this issue is not severe, especially when weighed against the benefit of reduced background efficiency, and the exact $N(b)=4$ selection is sufficient and effective., thereby selecting only events that reconstruct exactly four $b$-tagged jets.
+Following this pre-selection, we apply additional cuts on the transverse momenta of the leading and subleading $b$-jets, $P_{T}(b_{1})$ and $P_{T}(b_{2})$, for both BP1 and BP2 as both these kinematic variables show sufficient promise in isolating signal from background as can be seen from Figure 2.
+Figure 2: The plots in the upper panel represent the selection cuts applied to BP1.
 Both sets of plots show the selection cuts on $p_{T}(b_{1})$ and $p_{T}(b_{2})$.
@@
 However, even identifying the $4b$ channel and imposing a minimal set of cuts helps us isolate BSM effects quite clearly in this scenario.
-$BP-1$ : Cuts Signal (S) Background (B) S vs B
+BP1 : Cuts Signal (S) Background (B) S vs B
 Initial (no cut) 315 7509
@@
 SEL: $P_{T}(b_{2})>150$ 47.27 35.20 5.206
-$BP-2$ : Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 378 7509
-$N(b)=4$ 64.79 85.11 5.292
-SEL: $P_{T}(b_{1})>200$ 59.72 48.24 5.747
-SEL: $P_{T}(b_{2})>150$ 54.81 35.2 5.779
-Table 4: The upper table represents the cutflow chart for $BP-1$, while the lower table corresponds to $BP-2$.
-Both cases are considered at an integrated luminosity of $100\,\mathrm{fb}^{-1}$ in the analysis.
+Table 4: The table represents the cutflow chart for BP1.
+This is considered at an integrated luminosity of $100\,\mathrm{fb}^{-1}$ in the analysis.
 In this setup, a discovery potential of $5\sigma$ is anticipated for the signal.
@@
 This particular channel involving the production of $h_{1}H_{2}A_{2}$ and its subsequent decay into a fully hadronic mode with a $6b$ quark final state offers another promising avenue for exploration.
-The benchmark points considered here are referred to as $BP-3$ and $BP-4$ the corresponding masses, along with information regarding the branching ratios and total cross sections, are provided in Table 5.
+The benchmark point considered here is referred to as BP2 - the corresponding masses, along with information regarding the branching ratios and total cross sections, are provided in Table 5.
 A distinctive feature of this channel is its rare final state consisting of six $b$-quarks. The full production and decay chain is given below:
@@
 In Table 6, we list both irreducible and reducible backgrounds that are relevant for this channel along with their cross-sections.
-$BP-3$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $H_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
-Mass (GeV) 125 255.20 543.43 228.65 366.61 347.99 452.39 $\sigma||BR$ 0.0012179 0.54458 0.98325 0.981178 0.00063986
-$BP-4$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $h_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
-Mass (GeV) 125 269.54 470.15 215.39 522.60 311.28 451.15 $\sigma||BR$ 0.001619 0.54458 0.84123 0.82652 0.000613
-Table 5: The upper portion of the table represents the mass points, $BR$ and total cross section corresponding to $BP-3$, whereas the lower portion corresponds to $BP-4$.
-Background Process Final State Cross Section($\sigma$) (pb)
+BP2:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow h_{1}H_{2}A_{2}$ $h_{1}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ Total $\sigma$ (pb)
+Mass (GeV) 125 255.20 543.43 228.65 366.61 347.99 452.39 $\sigma\times BR$ 0.0012179 0.54458 0.98325 0.981178 0.00063986
+Table 5: The upper portion of the table represents the mass points, BR and total cross section corresponding to BP2.
+Background Process Final State Cross Section ($\sigma$) [pb]
 $e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}b\bar{b}$ $6b$ $5.886\times 10^{-6}$
@@
 The goal is to determine the range of luminosity required to effectively probe the production mechanism.
-The total cross sections for $BP-3$ and $BP-4$ are of similar magnitude, indicating that the required luminosity for discovery should also lie within a comparable range.
-We find that, for $BP-3$, an integrated luminosity of $1000\,\mathrm{fb}^{-1}$ and for $BP-4$, $1200\,\mathrm{fb}^{-1}$ are sufficient to achieve a discovery significance of $5\sigma$, based purely on the event count after applying the $N(b)=6$ selection.
+We find that, for BP2, an integrated luminosity of $1000\,\mathrm{fb}^{-1}$ are sufficient to achieve a discovery significance of $5\sigma$, based purely on the event count after applying the $N(b)=6$ selection.
 The corresponding selection cut and its effect on the event yield are summarized in the cutflow chart presented in Table 7.
-$BP-3$:Cuts Signal (S) Background (B) S vs B
+BP2 : Cuts Signal (S) Background (B) S vs B
 Initial (no cut) 639 92091
 SEL: $N(b)=6$ 30.71 5.0 5.14
-$BP-4$:Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 735 110509
-SEL: $N(b)=6$ 32.28 6.0 5.217
-Table 7: The upper portion of the table corresponds to the cutflow chart for $BP-3$ at $\mathcal{L}=1000\,fb^{-1}$, while the lower portion corresponds to $BP-4$ at $\mathcal{L}=1200\,fb^{-1}$.
+Table 7: The upper portion of the table corresponds to the cutflow chart for BP3 at $\mathcal{L}=1000\,fb^{-1}$, while the lower portion corresponds to BP4 at $\mathcal{L}=1200\,fb^{-1}$.
 III.4 $H_{2}H_{2}Z$ Production : $FHD$ Mode
@@
 Accordingly, a requirement of $N(b)\geq 4$ is imposed to retain signal efficiency while suppressing background contributions, and is adopted along with a cut of $N(j)\geq 2$.
-Following the structure of the previously discussed channels, we consider two benchmark points for the analysis, denoted as $BP\text{-}5$ and $BP\text{-}6$, which are consistent with current experimental constraints.
+Following the structure of the previously discussed channels, we consider a benchmark point for the analysis, denoted as BP3, which is consistent with current experimental constraints.
 The relevant SM backgrounds for this process are listed in Table 8.666This same background table is also referenced in the $A_{2}A_{2}Z$ production case, as both processes in the FHD mode yield the same final state of $4b+2j$.
 Detailed information on the chosen benchmark points—including their masses, branching ratios, and total production cross sections—is presented in Table 9.
-Background Process Final State Cross Section($\sigma$) (pb)
+Background Process Final State Cross Section ($\sigma$) [pb]
 $e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}jj$ $4b+2j$ 0.00141
@@
 $e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ 0.001098
-Table 8: The relevant SM backgrounds for the fully hadronic decay mode of $H_{2}H_{2}Z$ (Section III.4) and $A_{2}A_{2}Z$ (Section III.5) production.
-BP-5:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma||BR$ 0.007175 0.815618 0.815618 0.548 0.002616
-BP-6:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.548 0.002677
-Table 9: The upper table corresponds to $BP-5$, and the lower table to $BP-6$, summarizing the benchmark mass points along with the associated $BR$ and total cross sections.
+Table 8: The relevant SM backgrounds for the fully hadronic decay mode of $H_{2}H_{2}Z$ (subsection III.4) and $A_{2}A_{2}Z$ (subsection III.5) production.
+BP3:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
+Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma\times BR$ 0.007175 0.815618 0.815618 0.548 0.002616
+Table 9: The table corresponds to BP3, summarizing the benchmark mass points along with the associated BR and total cross sections.
 The selection cuts $N(b)\geq 4$ and $N(j)\geq 2$ are employed to preferentially isolate events consistent with the signal topology, thereby suppressing contributions from SM backgrounds.
 This approach, while simple in implementation, proves effective in enhancing the signal purity of the sample.
-Nevertheless, an important consequence of these inclusive number-based cuts is the significant overlap observed in the distributions of key kinematic observables between signal and background as can be evidenced from Fig.
-3.
+Nevertheless, an important consequence of these inclusive number-based cuts is the significant overlap observed in the distributions of key kinematic observables between signal and background as can be evidenced from Figure 3.
 As a result, no single observable emerges as a strong discriminator suitable for additional shape-based selection.
 Despite this limitation, the overall event yield remains sufficiently high.
-At an integrated luminosity of $300\,\mathrm{fb}^{-1}$, the signal reaches the threshold for discovery with a statistical significance of $5\sigma$.
+At an integrated luminosity of $500\,\mathrm{fb}^{-1}$, the signal reaches the threshold for discovery with a statistical significance of $5\sigma$.
 The cut-based selection criteria data is provided in Table 10.
 Figure 3: Some sample kinematic distributions for the $H_{2}H_{2}Z$ decaying to a fully hadronic final state - as can be seen, the overlap of the signal and background distributions for much of the kinematic range does not permit a useful cut-based strategy in this case.
-$BP-5$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 1000 30272
-SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 95.92 247.0 5.180
-$BP-6$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 1024 30272
-SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 112.9 247.0 5.951
-Table 10: The upper panel presents the cutflow summary for the BP-5 signal, while the lower panel shows the corresponding results for the BP-6 signal.
-Both cases assume an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
+BP3 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 1307 50454
+SEL: $N(b)\geq 4$ & $N(j)\geq 2$ 125.4 411.6 5.410
+Table 10: The panel presents the cutflow summary for the BP3 signal. Here we assume an integrated luminosity of $500\,\mathrm{fb}^{-1}$.
 III.5 $A_{2}A_{2}Z$ Production : $FHD$ Mode
-The $A_{2}A_{2}Z$ production channel and its subsequent decay topology are identical to the $H_{2}H_{2}Z$ case discussed in Section III.4.
+The $A_{2}A_{2}Z$ production channel and its subsequent decay topology are identical to the $H_{2}H_{2}Z$ case discussed in subsection III.4.
 The corresponding SM backgrounds are listed in Table 8.
-The signal benchmarks, labeled as $BP\text{-}7$ and $BP\text{-}8$, correspond to different mass configurations.
+The signal benchmarks, labeled as BP7 and BP8, correspond to different mass configurations.
 Detailed information on their mass values, BR’s and total production cross sections is provided in Table 11.
@@
 In particular, number-based selection criteria of $N(b)\geq 4$ and $N(j)\geq 2$ are applied, consistent with the methodology outlined in the previous section.
-The cutflow summary for benchmark points $BP\text{-}7$ and $BP\text{-}8$ is provided in Table 12.
+The cutflow summary for benchmark points BP4 is provided in Table 12.
 Notably, even with only these inclusive number cuts, a favorable balance between signal efficiency and background suppression is achieved, highlighting the effectiveness of this minimal selection strategy.
-BP-7:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.548 0.002341
-BP-8:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.548 0.002709
-Table 11: The upper table represents the mass points, BR and total cross section information for $BP-7$, whereas the lower table represents the same for $BP-8$.
-$BP-7$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 895 30272
-SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 108.43 247.0 5.752
-$BP-8$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 1036 30272
-SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 123.2 247.0 6.402
-Table 12: The upper panel presents the cutflow summary for the BP-7 signal, while the lower panel shows the corresponding results for the BP-8 signal.
-Both cases are analyzed at an integrated luminosity of $300\,\mathrm{fb}^{-1}$.
+BP4:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow jj$ Total $\sigma$ (pb)
+Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma\times BR$ 0.00637189 0.8188199 0.8188199 0.548 0.002341
+Table 11: The table represents the mass points, BR and total cross section information for BP4.
+BP4 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 936 40363
+SEL: $N(b)\geq 4$ & $N(j)\geq 2$ 113.42 329.3 5.391
+Table 12: The upper panel presents the cutflow summary for the BP4 signal. This case is analyzed at an integrated luminosity of $400\,\mathrm{fb}^{-1}$.
 III.6 $H_{2}H_{2}Z$ Production : $SLD$ Mode
-The $H_{2}H_{2}Z$ production channel discussed earlier in Section III.4 can also be considered with an alternative decay mode, wherein the $Z$ boson decays leptonically.
+The $H_{2}H_{2}Z$ production channel discussed earlier in subsection III.4 can also be considered with an alternative decay mode, wherein the $Z$ boson decays leptonically.
 The signal process is thus given by:
@@
 The relevant backgrounds are listed in Table 13777The background processes listed apply equally to both $H_{2}H_{2}Z$ and $A_{2}A_{2}Z$ production channels in the semi-leptonic decay mode, as they yield identical final states..
-Details of the two benchmark points, $BP\text{-}9$ and $BP\text{-}10$, including their masses, branching ratios, and cross sections, are provided in Table 14.
-Background Process Final State Cross Section($\sigma$) (pb)
+Details of the benchmark point, BP5, including their masses, branching ratios, and cross sections, are provided in Table 14.
+Background Process Final State Cross Section $\sigma$ (pb)
 $e^{+}e^{-}\rightarrow b\bar{b}b\bar{b}l^{+}l^{-}$ $4b+2l$ $1.893\times 10^{-5}$
@@
 […]
-Table 13: The relevant SM backgrounds for the semi-leptonic decay mode of $H_{2}H_{2}Z$ (Section III.6) and $A_{2}A_{2}Z$ (Section III.7) channels.
+Table 13: The relevant SM backgrounds for the semi-leptonic decay mode of $H_{2}H_{2}Z$ (subsection III.6) and $A_{2}A_{2}Z$ (subsection III.7) channels.
 The demand of at least two leptons necessitates the presence of SM gauge bosons that might also decay to neutrinos, and thus many of these backgrounds have sizable missing energy.
-BP-9:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
+BP5:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
 Mass (GeV) 125 207.83 527.73 315.17 397.45 333.29 417.01 $\sigma||BR$ 0.007175 0.815618 0.815618 0.0673 0.0003212
-BP-10:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}Z$ $H_{2}\rightarrow b\bar{b}$ $H_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
-Mass (GeV) 125 219.73 418.47 339.33 398.26 351.98 468.76 $\sigma||BR$ 0.006474 0.868647 0.868647 0.0673 0.0003288
-Table 14: The upper portion of the table represents the mass points, BR, and total cross-section information for $BP-9$, while the lower portion represents the same for $BP-10$.
+Table 14: The table represents the mass points, BR, and total cross-section information for BP5.
 The analysis begins by applying number-based selection criteria, requiring $N(b)\geq 4$ and $N(\ell)\geq 2$, to facilitate effective reconstruction of the signal.
@@
 Also given that the BR for leptonic decays is significantly lower than that of the hadronic mode discussed earlier, a higher integrated luminosity is generally required to achieve comparable signal sensitivity.
-The cutflow summary for benchmark points $BP\text{-}9$ and $BP\text{-}10$ is presented in Table 15.
-$BP-9$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 954 22147
-SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 40.85 17.04 5.368
-$BP-10$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 976 22147
-SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 49.23 17.04 6.05
-Table 15: The upper panel presents the cutflow summary for the BP-9 signal, while the lower panel shows the corresponding results for the BP-10 signal.
-Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
+The cutflow summary for benchmark points $BP5$ is presented in Table 15.
+BP5 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 963 33221
+SEL: $N(b)\geq 4$ &\& $N(\ell)\geq 2$ 41.26 25.6 5.047
+Table 15: The panel presents the cutflow summary for the BP5 signal. Here we assume an integrated luminosity of $3000\,\mathrm{fb}^{-1}$.
 III.7 $A_{2}A_{2}Z$ Production : $SLD$ Mode
-The $A_{2}A_{2}Z$ production channel followed by a semi-leptonic decay mode is analogous to the process discussed in Sec.
-III.6, yielding the same final state.
+The $A_{2}A_{2}Z$ production channel followed by a semi-leptonic decay mode is analogous to the process discussed in subsection III.6, yielding the same final state.
 Consequently, the relevant backgrounds remain unchanged and are listed in Table 13.
-Two benchmark mass points, denoted as $BP\text{-}11$ and $BP\text{-}12$, are considered, with their detailed properties provided in Table 16.
+A benchmark mass point, denoted as BP6, is considered, with its detailed properties provided in Table 16.
 The signal process is given by:
@@
 The resulting cutflow summary, following the implementation of these selection criteria, is presented in Table 17.
-BP-11:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
-Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma||BR$ 0.00637189 0.8188199 0.8188199 0.1 0.0002875
-BP-12:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
-Mass (GeV) 125 354.62 385.08 202.86 548.99 322.35 551.64 $\sigma||BR$ 0.00661869 0.864266 0.864266 0.0673 0.0003327
-Table 16: The upper portion of the table represents the mass points, BR, and total cross-section values for $BP-11$, while the lower portion represents the same for $BP-12$.
-$BP-11$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 853 22147
-SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 55.04 17.04 6.482
-$BP-12$: Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 988 22147
-SEL: $N(b)\geq 4$\&& $N(\ell)\geq 2$ 67.28 17.04 7.326
-Table 17: The upper panel presents the cutflow summary for the BP-11 signal, while the lower panel shows the corresponding results for the BP-12 signal.
-Both cases assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
+BP6:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}A_{2}Z$ $A_{2}\rightarrow b\bar{b}$ $A_{2}\rightarrow b\bar{b}$ $Z\rightarrow l^{+}l^{-}$ Total $\sigma$ (pb)
+Mass (GeV) 125 354.97 552.52 210.87 391.04 369.26 404.86 $\sigma\times BR$ 0.00637189 0.8188199 0.8188199 0.0673 0.0002875
+Table 16: The table represents the mass points, BR, and total cross-section values for BP6.
+BP5 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 575 22147
+SEL: $N(b)\geq 4$ &\& $N(\ell)\geq 2$ 37.07 17.04 5.04
+Table 17: The table presents the cutflow summary for the BP6 signal. Here we assume an integrated luminosity of $2000\,\mathrm{fb}^{-1}$.
 With these findings, we now move on to discussing more non-trivial production modes and their prospects, particularly in probing vertices involving the presence of charged Higgs bosons, as we continue our phenomenological discussions.
@@
 $e^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp},\quad(H_{2}\to b\bar{b}),\quad(H_{2}^{\pm}\to t\bar{b}/\bar{t}b\to W^{+}b\bar{b}/W^{-}b\bar{b}\to jjb\bar{b}),\quad(W^{\mp}\to jj)$
-Two representative benchmark points, $BP\text{-}13$ and $BP\text{-}14$, are considered in this study.
+A representative benchmark point, BP7 is considered in this study.
 Their mass configurations, BR and total production cross sections are summarized in Table 19.
@@
 The cutflow chart corresponding to this analysis, based on cut-based selections, is shown in Table 21.
-Background Process Final State Cross Section($\sigma$) (pb)
+Background Process Final State Cross Section $\sigma$ (pb)
 $e^{+}e^{-}\rightarrow ZZ\,b\bar{b}b\bar{b}\rightarrow jjjj\,b\bar{b}b\bar{b}$ $4b+4j$ $2.089\times 10^{-7}$
@@
 $e^{+}e^{-}\rightarrow t\bar{t}Z\rightarrow b\bar{b}b\bar{b}jjjj$ $4b+4j$ $2.774\times 10^{-4}$
-$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $0.06767$
-$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $0.02346$
-$e^{+}e^{-}\rightarrow t\bar{t}jj\rightarrow b\bar{b}jjjjjj$ $2b+6j$ $0.005821$
-$e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ $0.001098$
-Table 18: Relevant Standard Model backgrounds for the $H_{2}H_{2}^{\pm}W^{\mp}$ (Section III.8) and $A_{2}H_{2}^{\pm}W^{\mp}$ (Section III.9) production channels in the fully hadronic decay ($FHD$) mode, featuring a final state topology of $4b+4j$.
-$BP-13$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 224.16 359.96 377.98 474.54 326.93 500.67 $\sigma||BR$ 0.01095 0.8232 0.1575 1.0 0.6 0.001022
-$BP-14$:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 245.16 505.60 358.58 463.76 335.25 395.24 $\sigma||BR$ 0.01035 0.9159 0.1889 1.0 0.6 0.00129
-Table 19: The table provides information on $H_{2}H_{2}^{\pm}W^{\mp}$ production and its subsequent decay into a $4b+4j$ final state.
-The upper portion of the table corresponds to $BP-13$, while the lower portion corresponds to $BP-14$.
-$BP-13:$Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 1533 298496
-SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 104.0 302.8 5.15
-$BP-14:$Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 1934 298496
-SEL: $N(b)\geq 4$\&& $N(j)\geq 2$ 136.4 302.8 6.51
-Table 20: Cutflow charts for benchmark points $BP$-$13$ (upper table) and $BP$-$14$ (lower table). Both signals are analyzed at an integrated luminosity of $2000~{}\text{fb}^{-1}$.
+$e^{+}e^{-}\rightarrow t\bar{t}\rightarrow b\bar{b}jjjj$ $2b+4j$ $6.767\times 10^{-2}$
+$e^{+}e^{-}\rightarrow t\bar{t}j\rightarrow b\bar{b}jjjjj$ $2b+5j$ $2.346\times 10^{-2}$
+$e^{+}e^{-}\rightarrow t\bar{t}jj\rightarrow b\bar{b}jjjjjj$ $2b+6j$ $5.821\times 10^{-3}$
+$e^{+}e^{-}\rightarrow t\bar{t}jjj\rightarrow b\bar{b}jjjjjjj$ $2b+7j$ $1.098\times 10^{-3}$
+Table 18: Relevant Standard Model backgrounds for the $H_{2}H_{2}^{\pm}W^{\mp}$ (subsection III.8) and $A_{2}H_{2}^{\pm}W^{\mp}$ (subsection III.9) production channels in the fully hadronic decay (FHDFHD) mode, featuring a final state topology of $4b+4j$.
+BP7:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow H_{2}H_{2}^{\pm}W^{\mp}$ $H_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
+Mass (GeV) 125 224.16 359.96 377.98 474.54 326.93 500.67 $\sigma\times BR$ 0.01095 0.8232 0.1575 1.0 0.6 0.001022
+Table 19: The table provides information on $H_{2}H_{2}^{\pm}W^{\mp}$ production and its subsequent decay into a $4b+4j$ final state for BP7.
 III.9 $A_{2}H_{2}^{\pm}W^{\mp}$ Production : $FHD$ Mode
@@
 $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp},\quad(A_{2}\to b\bar{b}),\quad(H_{2}^{\pm}\to t\bar{b}/\bar{t}b\to W^{+}b\bar{b}/W^{-}b\bar{b}\to jjb\bar{b}),\quad(W^{\mp}\to jj)$
-We select two benchmark points—denoted as $BP\text{-}15$ and $BP\text{-}16$—that satisfy existing phenomenological constraints while offering representative signal kinematics.
+We select a benchmark point—denoted as BP8—that satisfy existing phenomenological constraints while offering representative signal kinematics.
 Their detailed properties, including masses, decay branching fractions, and production cross sections, are provided in Table 20.
@@
 Our results, summarized in the cutflow Table 22, indicate that with an integrated luminosity of $2000~\text{fb}^{-1}$, the signal can be probed with a discovery-level sensitivity of $5\sigma$.
-BP-15:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 365.69 370.25 222.74 366.90 309.51 437.50 $\sigma||BR$ 0.01111 0.96079 0.28799 1.0 0.6 0.002214
-BP-16:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
-Mass (GeV) 125 286.48 386.58 204.09 360.96 300.06 401.01 $\sigma||BR$ 0.011191 0.95317 0.24065 1.0 0.6 0.001848
-Table 21: The upper table presents the mass points, branching ratios ($BR$), and total cross section for $BP-15$, while the lower table provides the corresponding information for $BP-16$.
-$BP-15:$ Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 3320 298496
-SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 110.24 302.8 5.43
-$BP-16:$ Cuts Signal (S) Background (B) S vs B
-Initial (no cut) 2772 298496
-SEL: $N(b)\geq 4$\&& $N(j)\geq 4$ 143.73 302.8 6.8
-Table 22: The upper portion of the table presents the cutflow chart for $BP\text{-}15$, while the lower portion corresponds to $BP\text{-}16$. At an integrated luminosity of $2000~{}\text{fb}^{-1}$.
+BP8:Particle $h_{1}$ $H_{2}$ $H_{3}$ $A_{2}$ $A_{3}$ $H_{2}^{\pm}$ $H_{3}^{\pm}$ Channel $e^{+}e^{-}\rightarrow A_{2}H_{2}^{\pm}W^{\mp}$ $A_{2}\rightarrow b\bar{b}$ $H_{2}^{\pm}\rightarrow t\bar{b}/\bar{t}b$ $t/\bar{t}\rightarrow W^{+}b/W^{-}\bar{b}$ $W^{\pm}/W^{\mp}\rightarrow jj$ Total $\sigma$ (pb)
+Mass (GeV) 125 365.69 370.25 222.74 366.90 309.51 437.50 $\sigma\times BR$ 0.01111 0.96079 0.28799 1.0 0.6 0.002214
+Table 20: The table represents the mass points, branching ratios, and total cross-section values for BP8 for the $A_{2}H_{2}^{\pm}W^{\mp}\rightarrow 4b+4j$ final state.
+BP7 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 2043 198997
+SEL: $N(b)\geq 4$ &\& $N(j)\geq 4$ 104.0 302.8 5.15
+Table 21: Cut-flow results for benchmark points BP7. The signal and background events are analyzed at an integrated luminosity of $2000~\text{fb}^{-1}$.
+BP8 : Cuts Signal (S) Background (B) S vs B
+Initial (no cut) 4428 198997
+SEL: $N(b)\geq 4$ &\& $N(j)\geq 4$ 110.24 302.8 5.43
+Table 22: Cut-flow results for benchmark points BP8 at an integrated luminosity of $2000~\text{fb}^{-1}$.
 With all these channel surveys carried out, we now move on to the Discussion and Conclusion section section IV, where we highlight the relevant discussions and outline future outlooks that could be considered in experiments such as the ILC.
@@
 Our goal was to identify promising search strategies sensitive to interaction vertices characteristic of new physics, and to explore benchmark scenarios where viable production cross sections could be obtained, potentially leading to observable signals.
-We performed a detailed analysis of several production channels at the ILC, including $e^{+}e^{-}\to H_{2}A_{2}$, $e^{+}e^{-}\to H_{2}H_{2}Z$, $e^{+}e^{-}\to A_{2}A_{2}Z$, $e^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp}$, $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp}$, and $e^{+}e^{-}\to H_{1}H_{2}A_{2}$, which yield final states containing multiple jets, leptons, and b-quarks, making them ideal channels to investigate the extended Higgs sector.
+We performed a detailed analysis of several production channels at the ILC, including $^{+}e^{-}\to H_{2}A_{2}, e+e$$^{+}e^{-}\to H_{2}H_{2}Z$, $^{+}e^{-}\to A_{2}A_{2}Z, e+e$$^{+}e^{-}\to H_{2}H_{2}^{\pm}W^{\mp}$, $e^{+}e^{-}\to A_{2}H_{2}^{\pm}W^{\mp}$, and $e^{+}e^{-}\to H_{1}H_{2}A_{2}$, which yield final states containing multiple jets, leptons, and b-quarks, making them ideal channels to investigate the extended Higgs sector.
 For each channel, we selected representative benchmark points and conducted detailed phenomenological analysis.
@@
 However, the channels considered in our study do exhibit promising discovery potential at various luminosity thresholds.
-With the application of suitable number-based selection cuts and transverse momentum requirements on the leading and subleading $b$-quarks, the $H_{2}A_{2}$ channel yields a $5\sigma$ discovery significance at an integrated luminosity of $100~{}\text{fb}^{-1}$ for both benchmark points BP-1 and BP-2.
-For the $h_{1}H_{2}A_{2}$ channel, discovery is achievable at $1000~{}\text{fb}^{-1}$ for BP-3 and at $1200~{}\text{fb}^{-1}$ for BP-4.
-Similarly, the $H_{2}H_{2}Z$/$A_{2}A_{2}Z$ channel in the FHD mode requires approximately $300~{}\text{fb}^{-1}$ for discovery, whereas in the SLD mode, it requires about $2000~{}\text{fb}^{-1}$.
-Finally, the channel involving a charged Higgs, $H_{2}H_{2}^{\pm}W^{\mp}$/$A_{2}H_{2}^{\pm}W^{\mp}$, in the FHD mode demands a luminosity of up to $2000~{}\text{fb}^{-1}$ to reach discovery sensitivity.
-These results demonstrate that several production channels hinting at new physics can achieve a $5\sigma$ discovery at this collider setup.
+With the application of optimized selection cuts and transverse momentum requirements on the leading and subleading $b$-jets, the $H_{2}A_{2}$ channel achieves a $5\sigma$ discovery significance at an integrated luminosity of approximately $100~\text{fb}^{-1}$ for benchmark points BP1.
+For the $h_{1}H_{2}A_{2}$ channel, discovery is attainable at $1000~\text{fb}^{-1}$ for BP2.
+In the $H_{2}H_{2}A_{2}$ channel, the benchmark point BP3 require $500~\text{fb}^{-1}$, while BP5 need around $3000~\text{fb}^{-1}$ for discovery.
+The $A_{2}H_{2}A_{2}$ channel reaches $5\sigma$ at $400~\text{fb}^{-1}$ for BP4, and $2000~\text{fb}^{-1}$ for BP6.
+Finally, channels involving a charged Higgs boson, $H_{2}H_{2}^{\pm}W^{\mp}$ and $A_{2}H_{2}^{\pm}W^{\mp}$, demand an integrated luminosity close to $2000~\text{fb}^{-1}$ to achieve $5\sigma$ discovery significance for BP7 and BP8.
+These results collectively demonstrate that multiple new physics channels within this framework can reach discovery sensitivity at the considered collider setup.
 Overall, this study highlights promising directions for probing extended Higgs sectors at the ILC and providing the information of channel-wise analysis.
@@
 PMRF-192002-1802), awarded by the Ministry of Education, Government of India.
+We thank the anonymous referee for a careful reading of the manuscript and for several helpful comments and suggestions.